OpenStudy (anonymous):

Find the the slope of the line with the equation of 4x + 3y - 15=0 which is perpendicular.

OpenStudy (anonymous):

find the slope and then take the opposite reciprocal slope = -4/3 so perp line has a slope of 3/4

OpenStudy (sriram):

slope of given line=-4/3 ( y = -4x/3+15/3 ) slope of line Perpendicular=3/4 coz product of slopes=-1